\begin{thebibliography}{10}

\bibitem{Chau:1984fp}
L.-L. Chau and W.-Y. Keung, {\it {Comments on the Parametrization of the
  Kobayashi-Maskawa Matrix}},  {\em Phys.Rev.Lett.} {\bf 53} (1984) 1802.

\bibitem{Beringer:2012bj}
{\bf Particle Data Group} Collaboration, J.~Beringer et~al., {\it {The review
  of particle physics}},  {\em Phys. Rev.} {\bf D86} (2012) 010001.

\bibitem{Ahluwalia:2004dv}
D.~V. Ahluwalia-Khalilova, {\it {Addendum to: Gen. Rel. Grav. 28 (1996) 1161,
  First Prize Essay for 1996: Neutrino Oscillations and Supernovae}},  {\em
  Gen. Rel. Grav.} {\bf 36} (2004) 2183--2187.

\bibitem{Lunardini:2007vn}
C.~Lunardini, B.~Muller, and H.-T. Janka, {\it {Neutrino oscillation signatures
  of oxygen-neon-magnesium supernovae}},  {\em Phys. Rev.} {\bf D78} (2008)
  023016, [\href{http://xxx.lanl.gov/abs/0712.3000}{{\tt arXiv:0712.3000}}].

\bibitem{Duan:2006an}
H.~Duan, G.~M. Fuller, J.~Carlson, and Y.-Z. Qian, {\it {Simulation of Coherent
  Non-Linear Neutrino Flavor Transformation in the Supernova Environment. 1.
  Correlated Neutrino Trajectories}},  {\em Phys. Rev.} {\bf D74} (2006)
  105014, [\href{http://xxx.lanl.gov/abs/astro-ph/0606616}{{\tt
  astro-ph/0606616}}].

\bibitem{Duan:2007sh}
H.~Duan, G.~M. Fuller, J.~Carlson, and Y.-Z. Qian, {\it {Flavor Evolution of
  the Neutronization Neutrino Burst from an O-Ne-Mg Core-Collapse Supernova}},
  {\em Phys. Rev. Lett.} {\bf 100} (2008) 021101,
  [\href{http://xxx.lanl.gov/abs/0710.1271}{{\tt arXiv:0710.1271}}].

\bibitem{Yoshida:2006sk}
T.~Yoshida, T.~Kajino, H.~Yokomakura, K.~Kimura, A.~Takamura, et~al., {\it
  {Neutrino Oscillation Effects on Supernova Light Element Synthesis}},  {\em
  Astrophys. J.} {\bf 649} (2006) 319--331,
  [\href{http://xxx.lanl.gov/abs/astro-ph/0606042}{{\tt astro-ph/0606042}}].

\bibitem{Lesgourgues:2006nd}
J.~Lesgourgues and S.~Pastor, {\it {Massive neutrinos and cosmology}},  {\em
  Phys. Rept.} {\bf 429} (2006) 307--379,
  [\href{http://xxx.lanl.gov/abs/astro-ph/0603494}{{\tt astro-ph/0603494}}].

\bibitem{Khlopov:1981nq}
M.~Y. Khlopov and S.~Petcov, {\it {Possible cosmological effect of CP violation
  in neutrino oscillations}},  {\em Phys. Lett.} {\bf B99} (1981) 117.

\bibitem{Frampton:2002qc}
P.~Frampton, S.~Glashow, and T.~Yanagida, {\it {Cosmological sign of neutrino
  CP violation}},  {\em Phys. Lett.} {\bf B548} (2002) 119--121,
  [\href{http://xxx.lanl.gov/abs/hep-ph/0208157}{{\tt hep-ph/0208157}}].

\bibitem{Balantekin:2007es}
A.~B. Balantekin, J.~Gava, and C.~Volpe, {\it {Possible CP-Violation effects in
  core-collapse Supernovae}},  {\em Phys. Lett.} {\bf B662} (2008) 396--404,
  [\href{http://xxx.lanl.gov/abs/0710.3112}{{\tt arXiv:0710.3112}}].

\bibitem{Abe:2011ph}
{\bf Super-Kamiokande Collaboration} Collaboration, K.~Abe et~al., {\it {Search
  for Differences in Oscillation Parameters for Atmospheric Neutrinos and
  Antineutrinos at Super-Kamiokande}},  {\em Phys. Rev. Lett.} {\bf 107} (2011)
  241801, [\href{http://xxx.lanl.gov/abs/1109.1621}{{\tt arXiv:1109.1621}}].

\bibitem{Adamson:2011qu}
{\bf MINOS Collaboration} Collaboration, P.~Adamson et~al., {\it {Improved
  search for muon-neutrino to electron-neutrino oscillations in MINOS}},  {\em
  Phys. Rev. Lett.} {\bf 107} (2011) 181802,
  [\href{http://xxx.lanl.gov/abs/1108.0015}{{\tt arXiv:1108.0015}}].

\bibitem{Abe:2011fz}
{\bf DOUBLE-CHOOZ Collaboration} Collaboration, Y.~Abe et~al., {\it {Indication
  for the disappearance of reactor electron antineutrinos in the Double Chooz
  experiment}},  {\em Phys. Rev. Lett.} {\bf 108} (2012) 131801,
  [\href{http://xxx.lanl.gov/abs/1112.6353}{{\tt arXiv:1112.6353}}].

\bibitem{An:2012eh}
{\bf DAYA-BAY Collaboration} Collaboration, F.~An et~al., {\it {Observation of
  electron-antineutrino disappearance at Daya Bay}},  {\em Phys. Rev. Lett.}
  {\bf 108} (2012) 171803, [\href{http://xxx.lanl.gov/abs/1203.1669}{{\tt
  arXiv:1203.1669}}].

\bibitem{Ahn:2012nd}
{\bf RENO collaboration} Collaboration, J.~Ahn et~al., {\it {Observation of
  Reactor Electron Antineutrino Disappearance in the RENO Experiment}},  {\em
  Phys. Rev. Lett.} {\bf 108} (2012) 191802,
  [\href{http://xxx.lanl.gov/abs/1204.0626}{{\tt arXiv:1204.0626}}].

\bibitem{King:2012vj}
S.~King, {\it {Tri-bimaximal-Cabibbo Mixing}},
  \href{http://xxx.lanl.gov/abs/1205.0506}{{\tt arXiv:1205.0506}}.

\bibitem{Cabibbo:1963yz}
N.~Cabibbo, {\it {Unitary Symmetry and Leptonic Decays}},  {\em Phys. Rev.
  Lett.} {\bf 10} (1963) 531--533.

\bibitem{Wolfenstein:1983yz}
L.~Wolfenstein, {\it {Parametrization of the Kobayashi-Maskawa Matrix}},  {\em
  Phys. Rev. Lett.} {\bf 51} (1983) 1945.

\bibitem{Mohapatra:2006gs}
R.~Mohapatra and A.~Smirnov, {\it {Neutrino Mass and New Physics}},  {\em Ann.
  Rev. Nucl. Part. Sci.} {\bf 56} (2006) 569--628,
  [\href{http://xxx.lanl.gov/abs/hep-ph/0603118}{{\tt hep-ph/0603118}}].

\bibitem{Fukuda:1998mi}
{\bf Super-Kamiokande Collaboration} Collaboration, Y.~Fukuda et~al., {\it
  {Evidence for oscillation of atmospheric neutrinos}},  {\em Phys. Rev. Lett.}
  {\bf 81} (1998) 1562--1567,
  [\href{http://xxx.lanl.gov/abs/hep-ex/9807003}{{\tt hep-ex/9807003}}].

\bibitem{Ahluwalia:2002tr}
D.~V. Ahluwalia, Y.~Liu, and I.~Stancu, {\it {CP-violation in neutrino
  oscillations and $L/E$ flatness of the E-like event ratio at
  Super-Kamiokande}},  {\em Mod. Phys. Lett.} {\bf A17} (2002) 13--21.

\bibitem{Zhang:2012ys}
X.~Zhang and B.-Q. Ma, {\it {A Prediction of neutrino mixing matrix with CP
  violating phase}},  \href{http://xxx.lanl.gov/abs/1203.2906}{{\tt
  arXiv:1203.2906}}.

\bibitem{Tortola:2012te}
M.~Tortola, J.~Valle, and D.~Vanegas, {\it {Global status of neutrino
  oscillation parameters after recent reactor measurements}},
  \href{http://xxx.lanl.gov/abs/1205.4018}{{\tt arXiv:1205.4018}}.

\bibitem{Fogli:2012ua}
G.~Fogli, E.~Lisi, A.~Marrone, D.~Montanino, A.~Palazzo, et~al., {\it {Global
  analysis of neutrino masses, mixings and phases: entering the era of leptonic
  CP violation searches}},  \href{http://xxx.lanl.gov/abs/1205.5254}{{\tt
  arXiv:1205.5254}}.

\bibitem{Harrison:2002er}
P.~Harrison, D.~Perkins, and W.~Scott, {\it {Tri-bimaximal mixing and the
  neutrino oscillation data}},  {\em Phys. Lett.} {\bf B530} (2002) 167,
  [\href{http://xxx.lanl.gov/abs/hep-ph/0202074}{{\tt hep-ph/0202074}}].

\bibitem{Stancu:1999ct}
I.~Stancu and D.~V. Ahluwalia, {\it {L / E flatness of the electron - like
  event ratio in Super-Kamiokande and a degeneracy in neutrino masses}},  {\em
  Phys. Lett.} {\bf B460} (1999) 431--436,
  [\href{http://xxx.lanl.gov/abs/hep-ph/9903408}{{\tt hep-ph/9903408}}].

\bibitem{Vissani:1997pa}
F.~Vissani, {\it {A Study of the scenario with nearly degenerate Majorana
  neutrinos}},  \href{http://xxx.lanl.gov/abs/hep-ph/9708483}{{\tt
  hep-ph/9708483}}.

\bibitem{Ahluwalia:1998xb}
D.~V. Ahluwalia, {\it {Reconciling Super-Kamiokande, LSND, and home-stake
  neutrino oscillation data}},  {\em Mod. Phys. Lett.} {\bf A13} (1998)
  2249--2264, [\href{http://xxx.lanl.gov/abs/hep-ph/9807267}{{\tt
  hep-ph/9807267}}].

\bibitem{Barger:1998ta}
V.~D. Barger, S.~Pakvasa, T.~J. Weiler, and K.~Whisnant, {\it {Bimaximal mixing
  of three neutrinos}},  {\em Phys. Lett.} {\bf B437} (1998) 107--116,
  [\href{http://xxx.lanl.gov/abs/hep-ph/9806387}{{\tt hep-ph/9806387}}].

\bibitem{Wolfenstein:1977ue}
L.~Wolfenstein, {\it {Neutrino Oscillations in Matter}},  {\em Phys.Rev.} {\bf
  D17} (1978) 2369--2374.

\bibitem{Mikheev:1986gs}
S.~Mikheev and A.~Y. Smirnov, {\it {Resonance Amplification of Oscillations in
  Matter and Spectroscopy of Solar Neutrinos}},  {\em Sov. J. Nucl. Phys.} {\bf
  42} (1985) 913--917.

\bibitem{Gava:2008rp}
J.~Gava and C.~Volpe, {\it {Collective neutrinos oscillation in matter and
  CP-violation}},  {\em Phys. Rev.} {\bf D78} (2008) 083007,
  [\href{http://xxx.lanl.gov/abs/0807.3418}{{\tt arXiv:0807.3418}}].

\bibitem{Kneller:2009vd}
J.~P. Kneller and G.~C. McLaughlin, {\it {Three Flavor Neutrino Oscillations in
  Matter: Flavor Diagonal Potentials, the Adiabatic Basis and the CP phase}},
  {\em Phys.Rev.} {\bf D80} (2009) 053002,
  [\href{http://xxx.lanl.gov/abs/0904.3823}{{\tt arXiv:0904.3823}}].

\bibitem{Kisslinger:2012se}
L.~S. Kisslinger, E.~M. Henley, and M.~B. Johnson, {\it {Neutrino Oscillation
  in Matter and Parameters $s_{13},\delta_{CP}$}},
  \href{http://xxx.lanl.gov/abs/1203.6613}{{\tt arXiv:1203.6613}}.

\bibitem{Ahluwalia:1996ev}
D.~V. Ahluwalia and C.~Burgard, {\it {Gravitationally induced quantum
  mechanical phases and neutrino oscillations in astrophysical environments}},
  {\em Gen. Rel. Grav.} {\bf 28} (1996) 1161--1170,
  [\href{http://xxx.lanl.gov/abs/gr-qc/9603008}{{\tt gr-qc/9603008}}].

\bibitem{Ahluwalia:1998jx}
D.~V. Ahluwalia and C.~Burgard, {\it {Interplay of gravitation and linear
  superposition of different mass eigenstates}},  {\em Phys. Rev.} {\bf D57}
  (1998) 4724--4727, [\href{http://xxx.lanl.gov/abs/gr-qc/9803013}{{\tt
  gr-qc/9803013}}].

\bibitem{Konno:1998kq}
K.~Konno and M.~Kasai, {\it {General relativistic effects of gravity in quantum
  mechanics: A Case of ultrarelativistic, spin 1/2 particles}},  {\em Prog.
  Theor. Phys.} {\bf 100} (1998) 1145--1157,
  [\href{http://xxx.lanl.gov/abs/gr-qc/0603035}{{\tt gr-qc/0603035}}].

\bibitem{Wudka:2000rf}
J.~Wudka, {\it {Mass dependence of the gravitationally induced wave function
  phase}},  {\em Phys. Rev.} {\bf D64} (2001) 065009,
  [\href{http://xxx.lanl.gov/abs/gr-qc/0010077}{{\tt gr-qc/0010077}}].

\bibitem{Mukhopadhyay:2005gb}
B.~Mukhopadhyay, {\it {Neutrino asymmetry around black holes: Neutrinos
  interact with gravity}},  {\em Mod. Phys. Lett.} {\bf A20} (2005) 2145--2156,
  [\href{http://xxx.lanl.gov/abs/astro-ph/0505460}{{\tt astro-ph/0505460}}].

\bibitem{Singh:2003sp}
P.~Singh and B.~Mukhopadhyay, {\it {Gravitationally induced neutrino
  asymmetry}},  {\em Mod. Phys. Lett.} {\bf A18} (2003) 779--785.
\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\pdfoutput=1
\usepackage{jheppub}

\usepackage{amsmath,colonequals}
\usepackage{amssymb}
\usepackage{graphicx}
\usepackage{epsfig}
\usepackage{bm}
\usepackage{color}

\newfont{\sss}{cmmi10 at 20.74pt}

\def\spose#1{\hbox to 0pt{#1\hss}}
\def\simlt{\mathrel{\spose{\lower 3pt\hbox{$\mathchar"218$}}
     \raise 2.0pt\hbox{$\mathchar"13C$}}}
\def\simgt{\mathrel{\spose{\lower 3pt\hbox{$\mathchar"218$}}
     \raise 2.0pt\hbox{$\mathchar"13E$}}}
\def\0{\mbox{\boldmath$\displaystyle\mathbf{0}$}}

\def\be{\begin{equation}}
\def\ee{\end{equation}}
\def\bea{\begin{eqnarray}}
\def\eea{\end{eqnarray}}

\newcommand{\ce}{\colonequals}

\title{CP violatingTri-bimaximal-Cabibbo mixing}

\author{D. V. Ahluwalia}

\affiliation{Department of Physics and Astronomy,
Rutherford Building, University of Canterbury,
Private Bag 4800, Christchurch 8140, New Zealand}

\emailAdd{dharamvir.ahluwalia@canterbury.ac.nz}

\keywords{Neutrino physics, CP violation}

\abstract{ In view of the new data from the Daya Bay and RENO collaborations, King has presented a very natural deformation of tri-bimaximal mixing. Here we show that $L/E$ flatness of the $e$-like event ratio in the atmospheric neutrino data, when coupled with King's observation that the smallest neutrino mixing angle, $\theta_{13}$, seems to be related to the largest 
quark mixing angle (the Cabibbo angle $\theta_C$), leads to a CP violating  tri-bimaximal-Cabibbo mixing. King's tri-bimaximal-Cabibbo mixing follows as a leading order approximation from our result.  }

\notoc
\begin{document}

\maketitle
\bigskip\bigskip

\hrule
The precise form of the neutrino mixing matrix, $U$,
that defines the relationship between the flavour and mass eigenstates, $\vert\nu_\ell\rangle$ and $\vert\nu_j\rangle$ respectively~\cite{Chau:1984fp,Beringer:2012bj}, reads 
\begin{equation}
\vert\nu_\ell\rangle = \sum_j U^\ast_{\ell j} \vert\nu_j\rangle, \quad\ell = e,\mu,\tau,\quad j=1,2,3,
\end{equation}
and  the knowledge of the masses for the underlying mass eigenstates, arise from yet unknown physics.  Nevertheless, once the parameters that determine the mixing matrix and the mass-squared differences are deciphered from the data one can derive their phenomenological consequences on supernova explosions~\cite{Ahluwalia:2004dv,Lunardini:2007vn,Duan:2006an,Duan:2007sh},  on the synthesis of elements~\cite{Yoshida:2006sk},  on the cosmic microwave background and the distribution of large-scale structure~\cite{Lesgourgues:2006nd}. In particular, if the neutrino mixing angle $\theta_{13} \ne 0$ then one can obtain CP violation in the neutrino sector with many interesting physical consequences~\cite{Khlopov:1981nq,Frampton:2002qc,Balantekin:2007es}.

The T2K, MINOS, and Double CHOOZ indications that the smallest neutrino mixing angle $\theta_{13}$ may  be non-zero~\cite{Abe:2011ph,Adamson:2011qu,Abe:2011fz} has now been confirmed by the results of the Daya Bay and RENO collaborations~\cite{An:2012eh,Ahn:2012nd}. King has  made the observation~\cite{King:2012vj} that the smallest neutrino mixing angle $\theta_{13}$, seems to be related to the largest 
quark mixing angle, the Cabibbo angle $\theta_C$~\cite{Cabibbo:1963yz}, or equivalently to the Wolfenstein parameter, $\lambda = 0.2253\pm0.0007$~\cite{Wolfenstein:1983yz,Beringer:2012bj}:\footnote{It is worth noting that Mohapatra and Smirnov had earlier conjectured King's observation~\cite[Sec.~3.1]{Mohapatra:2006gs}.}
\begin{equation}
 \theta_{13} ~ \mbox{(or, } \theta_{reac}\mbox{)} = \arcsin\left(\frac{\sin\theta_C}{\sqrt{2}}\right) = \arcsin\left(\frac{\lambda}{\sqrt{2}}\right).
\end{equation}
To this observation we now add that the $L/E$ \textemdash~where $L$ is the neutrino source-detector distance and $E$ is the neutrino energy ~\textemdash ~flatness of the $e$-like event ratio observed for atmospheric neutrinos~\cite{Fukuda:1998mi} requires that
\begin{equation}
\theta_{23}~ \mbox{(or, } \theta_{atm}\mbox{)}  = \frac{\pi}{4}, \quad \delta= \pm \frac{\pi}{2}.\label{eq:2}
\end{equation}
This observation was first made in reference~\cite{Ahluwalia:2002tr}.  The $\delta$ obtained in~\cite{Ahluwalia:2002tr} was also introduced recently as an Ansatz in Ref.~\cite{Zhang:2012ys}.

Global analysis of neutrino oscillation data by two independent groups shows: (a) $\delta$ to be $\left(0.83^{+0.54}_{-0.64}\right)\pi$ for the normal mass hierarchy while allowing for the full $[0,2 \pi]$ range for the inverted mass hierarchy~\cite{Tortola:2012te}, (b) $\delta \approx \pi$ with no significant difference between the normal and inverted mass hierarchies~\cite{Fogli:2012ua}. A detailed study of these two papers reveals that there is no statistically significant indication which disfavours $\delta  = \pm \pi/2$. Regarding $\theta_{23}$: (a)  the first of the mentioned groups obtains $\sin^2\theta_{23} = 0.49^{+0.08}_{-0.05}$ for the normal mass hierarchy, and $\sin^2\theta_{23} = 0.53^{+0.05}_{-0.07}$ for the inverted mass hierarchy (these values are consistent with $\theta_{23}=\pi/4$), while (b) the second group finds a slight preference for $\theta_{23} < \pi/4$.

Both groups agree with the tri-bimaximal mixing value for the remaining angle~\cite{Tortola:2012te,Fogli:2012ua}
\begin{equation}
\theta_{12}~ \mbox{(or, } \theta_{\odot}\mbox{)}  =  \arcsin\left(\frac{1}{\sqrt{3}}\right).
\end{equation}
With all the angles and phases thus fixed, the neutrino mixing matrix for the choice $\delta  = \pi/2$ in equation~(\ref{eq:2}) takes the form

\begin{equation}
U^+ = \begin{pmatrix}
\sqrt{\frac{2}{3}}
\left(1-\frac{\lambda^2}{2}\right)^{1/2} & 
\sqrt{\frac{1}{3}} \left(1-\frac{\lambda^2}{2}\right)^{1/2} & i \frac{1}{\sqrt{2}}\lambda\\ 
-\frac{1}{\sqrt{6}} \left(1- i \lambda\right) &
\frac{1}{\sqrt{3}}\left(1 + i \frac{1}{2}\lambda\right) &
\frac{1}{\sqrt{2}}\left( 1-\frac{\lambda^2}{2} \right)^{1/2}\\
\frac{1}{\sqrt{6}} \left(1+ i \lambda\right) &
-\frac{1}{\sqrt{3}}\left(1 - i \frac{1}{2}\lambda\right) &
\frac{1}{\sqrt{2}}\left( 1-\frac{\lambda^2}{2} \right)^{1/2}
\end{pmatrix}.
\end{equation}
Its counterpart, $U^-$, for $\delta = - \pi/2$ is obtained by letting $i\to -i$ in $U^+$. As a measure of CP violation, following~\cite{Beringer:2012bj}, we define the asymmetries
\begin{equation}
A_{CP}^{(\ell^\prime\ell)}\ce P(\nu_\ell \to \nu_{\ell^\prime}) -
 P(\bar\nu_\ell \to \bar\nu_{\ell^\prime}),
\end{equation}
and find 
\begin{align}
A_{CP}^{(\mu e)} = - A^{(\tau e)}_{CP} = A_{CP}^{(\tau\mu)} & =  \mp\frac{1}{3} \lambda\left(2 - \lambda^2\right) \left(
\sin \frac{\Delta m^2_{32}}{ 2 p} L
+ \sin \frac{\Delta m^2_{21}}{ 2 p} L
+ \sin \frac{\Delta m^2_{13}}{ 2 p} L
\right)\nonumber\\
&\approx \mp0.146 \left(
\sin \frac{\Delta m^2_{32}}{ 2 p} L
+ \sin \frac{\Delta m^2_{21}}{ 2 p} L
+ \sin \frac{\Delta m^2_{13}}{ 2 p} L
\right),\label{eq:cpv}
\end{align}
where all symbols have their usual meaning. The $\mp$ sign holds for $\delta =\pm \frac{\pi}{2}$. For $\lambda = 0$, or equivalently $\theta_{13}=0$, the $U^\pm$ reduce to the standard tri-bimaximal mixing 
matrix~\cite{Harrison:2002er}.\footnote{This may be compared with~\cite[Eq.~(26)]{Stancu:1999ct} that gives an interpolating matrix with $\theta_\odot$ as a variable. In one limit the interpolating matrix gives the bimaximal mixing~\cite{Vissani:1997pa,Ahluwalia:1998xb,Barger:1998ta} and in another it yields tri-bimaximal mixing~\cite{Harrison:2002er}. }

The result~(\ref{eq:cpv}) is modified by matter effects~\cite{Wolfenstein:1977ue,Mikheev:1986gs}. Its general features are studied in detail by various authors~\cite{Gava:2008rp,Balantekin:2007es,Kneller:2009vd,Kisslinger:2012se}.
In gravitational environments the following argument suggests that one must expect a significant modification to the result~(\ref{eq:cpv}). Neutrino oscillations provide us with a set of flavour oscillation clocks. These clocks must redshift according to the general expectations of the theory of general relativity.
In gravitational environments of neutron stars the dimensionless gravitational potential is $\Phi^{NS}_{grav}\approx 0.2$ (cf. for Earth, $\Phi^{\oplus}_{grav}\approx 6.95 \times 10^{-10}$). For a given source-detector distance, and a given energy, the asymmetries $A_{CP}$ for supernovae modeling must be accordingly modified~\cite{Ahluwalia:1996ev,Ahluwalia:1998jx,Konno:1998kq,Wudka:2000rf,Mukhopadhyay:2005gb,Singh:2003sp} at the $20\%$ level, or thereabouts.

 An examination of the $U^\pm$ immediately shows that the expectation values of the $\nu_\mu$ and $\nu_\tau$ masses are identical.
To $\mathcal{O}(\lambda^2)$ the $U^-$ obtained above reproduces to King's result~\cite[Eq.~(8)]{King:2012vj} for $\delta = \pi/2$.
The presented $U^\pm$ not only accommodate the implications of the Daya Bay and RENO collaborations, but also the L/E flatness of the $e$-like event ratio seen in the atmospheric neutrino data while respecting all other known data on neutrino oscillations.

\acknowledgments
The result presented here was obtained on 10 May 2012, and was presented the next day at a MatScience Seminar. The author thanks 
Institute of Mathematical Sciences (``MatScience'', Chennai, India)   for its hospitality and for its vibrant scholarly environment.


\providecommand{\href}[2]{#2}\begingroup\raggedright
\endgroup

\end{document}